OpenStudy (victoriathebear):

Simplify and solve this equation: 4m + 9 + 5m –12 = 42. A. m = -4 (B.) m = 5 C. m = 4 D. m = –5 could you explain if I'm correct?

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 4m + 9 + 5m –12 = 42 }\) \(\normalsize\color{blue}{ 4m + 5m –12+ 9 = 42 }\) \(\normalsize\color{blue}{ 9m –3 = 42 }\) \(\normalsize\color{blue}{ 9m –3 \color{red}{+3} = 42 \color{red}{+3} }\) \(\normalsize\color{blue}{ 9m = 45 }\) \(\normalsize\color{blue}{ 9m \color{red}{\div 9} = 45\color{red}{\div 9} }\) \(\normalsize\color{blue}{ m = 5 }\)

OpenStudy (anonymous):

so you gave her the answer smh

OpenStudy (yanasidlinskiy):

He still showed her how to do it though:D She can now work out similar problems just like this:D

OpenStudy (victoriathebear):

i did it myself thankyou @sunrise14 i was just checking if I was correct.

OpenStudy (solomonzelman):

what is `smh` supposed to stay for, by the way ?

OpenStudy (victoriathebear):

shake my head -.- thankyou @SolomonZelman

OpenStudy (yanasidlinskiy):

Lol:D He's the smartest:D

OpenStudy (yanasidlinskiy):

I caught ya:D It's *stand* for. Not "stay for"!!!:D Lol. I'm just Kidding:D

OpenStudy (victoriathebear):

:D

OpenStudy (yanasidlinskiy):

If i got $1 for every mistake I did......I would be flowing with the money!!:D I would be a millionaire by now:D

OpenStudy (solomonzelman):

Everyone one would be a millionaire :) Certainly me , no joke.

OpenStudy (yanasidlinskiy):

Hahhaha!!:D Solomon!!:D Especially you!!:D I caught you 2wice maybe already!!:D